what do we find actually in disatance travelled in nth second

Grade:11

1 Answers

Shaswata Biswas
7 years ago
We have distance travellen in t second, S = u.t + a.t2/2
Distance travelled in tth = ST - ST-1 
= {u.t + a.t2/2} - {u.(t-1) + a.(t-1)2/2}
= u{t - (t-1)} + a{t2 - (t-1)2}/2
= u + a(2t-1)/2
 
So the distance travelled in nth second, Sn = u + a(2n-1)/2
THANKS